Use a contour to evaluate So S**** dx T 4 8 Oy O | EIS 2
Answer #1

lleg I= x3 abe ti it is a simple calculas Questions let it so nit x3 lim I = dra to x 8+1 Put xt = u => 4x² cha=du x3 dx=du 4I I lim 4 t 50 tant (at) - tant cos (tail (oo) tant (60) - o (teh ( tor(E.) I = 1 x -H H lt I = 2 I = - 5lt